# Data-Efficient Operator Learning via Unsupervised Pretraining and In-Context Learning

### v2

@@ 5 Conclusion
In this work, we focus on enhancing the data efficiency of solving PDEs using deep networks, particularly through unsupervised pretraining and ICL methods. Our key contributions include the introduction of unsupervised pretraining for operator learning and a flexible ICL approach that enhances OOD generalization without incurring extra training costs. Through extensive evaluations, we demonstrate that our method is not only data-efficient but also more generalizable compared to existing approaches. We anticipate that our work would inspire the SciML community to explore more on solving the heavy simulation costs and unsatisfactory OOD generalization of neural operators. Current limitations include: 1) more proxy tasks could be further designed; 2) larger SciML foundation models could be studied; 3) more physical constraints (e.g. the conservation law) could be considered. These limitations are beyond the scope of this work, and we expect to address them in future research.
Acknowledgements
This work was supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Scientific Discovery through Advanced Computing (SciDAC) program, under Contract Number DE-AC02-05CH11231 at Lawrence Berkeley National Laboratory. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility located at Lawrence Berkeley National Laboratory, operated under Contract Number DE-AC02-05CH11231 using NERSC award ASCR-ERCAP0023337.
@@ Appendix A PDEs
    $\begin{split}\partial_{t}w(x,t)+u(x,t)\cdot\nabla w(x,t)&=\nu\Delta w(x,t)+f(x),\\ \nabla\cdot u(x,t)&=0,\\ w(x,0)&=w_{0}(x).\end{split}$ (5)
    Here, the velocity $u$ and the vorticity $w$ are defined within the time duration $[0,T]$ and the spatial domain $[0,1]^{2}$ (the vorticity can also be formulated as $w=\nabla\times u$), and $w_{0}$ is the initial vorticity. Moreover, the coefficient $\nu$ signifies the viscosity of a fluid, and $f(x)$ is the external forcing function. Periodic boundary conditions are employed in this system.
    Unlabeled PDE Data: We follow the training settings from previous works. 1) For FNO [4], we focus on mapping the vorticity up to time $T$ to the vorticity at a later time $T+1$ ($\left.w_{[0,T]}\mapsto w\right|_{T+1}$) (with $T=33$). The input includes four channels: the vorticity, and three (xy-spatial and temporal) linear mesh position embedding channels. Thus, similar to Reaction-Diffusion above, during pretraining, the model will take individual snapshots of vorticity and position embeddings as unlabeled data. 2) For PDEBench [10], we forecast both xy velocities and vorticity (i.e. three input channels), with $T=15$. The model will take individual snapshots of vorticity and xy velocities as unlabeled data.
@@ B.1 Distributions of Unlabeled PDE Data
In Table 1, we summarize the distribution of our unlabeled PDE data during pretraining, fine-tuning, and OOD inference with ICL. Ranges of these physical parameters are inspired by [11]. During pretraining, we consider a wide distribution of unlabeled PDE data. When training (fine-tuning) our model, we consider in-distribution unlabeled PDE data. Finally, we test our ICL method with OOD settings.333 For Navier Stokes from PDEBench, we use the original data. We could not generate our own pretraining/finetuning data with different ranges of physics parameters, due to a possible mismatch of the provided configuration files and the version of Phiflow used in PDEBench (see GitHub issue at https://github.com/pdebench/PDEBench/issues/36). For Helmholtz OOD, we choose a narrow range of coefficients ([15, 20]) mainly because its solution is very sensitive to the wavenumber, and FNO’s performance significantly drops when we move to more extreme OOD settings.
Table 1: Ranges of physical parameters (integers) for unsupervised pretraining, training (fine-tuning), and out-of-distribution (OOD, for ICL).
@@ B.3 Training Hyperparameters
64$\times$64 64$\times$64 128$\times$128$\times$10 64$\times$64$\times$1 512$\times$512$\times$15 180$\times$180$\times$16 180$\times$120$\times$16 96$\times$45
@@
Rollouts N/A N/A 91 N/A 1 N/A N/A N/A
Appendix C Example of Simulation Costs
@@ Appendix D Model Architectures
In Figure 7, We show visualizations of architectures described in Sec. 3.1.4.
@@ Appendix E Comparison with Contrastive Learning
Appendix F Consistently Improved Generalization
@@
Appendix G More Ablation Studies
G.1 Magnitude of Perturbations during Pretraining.
@@ Appendix G More Ablation Studies
G.2 Ablation of the Number of Pretraining Samples.
@@
Appendix H Visualization of MAE Pretraining
@@
Appendix I Benefits of In-context Learning (ICL)
@@
We try to understand the benefit of ICL by decomposing the relative MSE error into “scale” and “shape.” “Scale” is the slope of a linear regression between targets and predictions (closer to 1 the better), indicating the alignment of the range of model outputs with targets. “Shape” is the normalized relative MSE (i.e., model outputs or targets are normalized by their own largest magnitude before MSE), indicating the alignment of scale-invariant spatial/temporal structures. We show results in Figure 12. We find that the benefit of ICL lies in that the scale of the model’s output keeps being better calibrated (“scale” being closer to 1) when adding more demos.
@@
Figure 12: Benefits of ICL. To analyze the benefit of ICL in complicated PDE systems, we decompose the relative MSE error into “Scale” and “Shape”. “Scale” indicates the alignment of the range of model outputs with targets (closer to 1 the better), via the slope of a linear regression. “Shape” indicates the alignment of scale-invariant spatial/temporal structures via normalized relative MSE (i.e. model outputs or targets are normalized by their own largest magnitude before MSE). We find that the benefit of ICL lies in that the scale of the model’s output keeps being calibrated (red line being closer to 1) when adding more demos.

### v3

@@ 5 Conclusion
In this work, we focus on improving the data efficiency of solving partial differential equations (PDEs) using deep learning, with a particular emphasis on unsupervised pretraining and in-context learning (ICL) methods. Our key contributions include introducing unsupervised pretraining for operator learning and a flexible ICL approach that enhances out-of-distribution (OOD) generalization without increasing training costs. Through extensive evaluations, we demonstrate that our method is not only more data-efficient, but it also achieves greater generalizability compared to existing approaches. By improving the data efficiency of neural operators for solving PDEs, our approach can significantly reduce the computational costs and energy demands of high-fidelity numerical PDE simulations. Additionally, by making advanced PDE solutions more accessible through efficient pretraining, our method has the potential to accelerate scientific and engineering progress across various fields, ultimately benefiting society. We hope our work will inspire the scientific machine learning (SciML) community to further address the high simulation costs and limited OOD generalization of neural operators, contributing to advancements that support both scientific innovation and environmental sustainability.
6 Limitations
Current limitations of our work: 1) We could design more physics-inspired proxy tasks and data augmentation methods for scientific data; 2) We could study more PDE systems in our unsupervised pretraining and in-context learning; 3) We could consider more different neural operator architectures. We expect that addressing these limitations will lead to broader impacts in future works.
@@ Appendix A PDEs
    $\begin{split}\nabla\cdot\mathbf{v}&=0\\ \rho\left(\partial_{t}\mathbf{v}+\mathbf{v}\cdot\nabla\mathbf{v}\right)&=-\nabla p+\nu\nabla^{2}\mathbf{v}+\mathbf{f}\end{split}$ (5)
    Here, the velocity $\mathbf{v}$ is defined within the time duration $[0,T]$ and the spatial domain $[0,1]^{2}$ (the vorticity can be formulated as $w=\nabla\times\mathbf{v}$). Moreover, $\rho$ is the density, and the coefficient $\nu$ signifies the viscosity of a fluid, and $\mathbf{f}$ is the external forcing function. Dirichlet boundary conditions are employed in this system.
    Unlabeled PDE Data: We follow the training settings from previous works. 1) For FNO [4], we focus on mapping the vorticity from the initial condition to the full solution ($\left.w_{0}\mapsto w\right|_{[0,T)}$, with $T=33$). The input includes four channels: the vorticity (which is the initial condition, duplicated for $T$ times), and three (xy-spatial and temporal) linear mesh position embedding channels. Thus, similar to Reaction-Diffusion above, during pretraining, the model will take individual snapshots of vorticity and position embeddings as unlabeled data. 2) For PDEBench [10], we forecast both xy velocities and pressure (i.e. three input channels), with $T=15$. The model will take individual snapshots of vorticity and xy velocities as unlabeled data.
We summarize detailed inputs and outputs in Table 1.
Table 1: Inputs and outputs for learning different PDEs. See Table 3 for resolutions. “NS”: Navier-Stokes. “RD”: Reaction-Diffusion.
PDE Simulations Input Input Shape Output
Poisson [11] Source function ($f$), diffusion coefficients ($\bm{K}$) $C\times H\times W(C=4)$ Potential field ($u$)
Helmholtz [11] Source function ($f$), wavenumber $\omega$ $C\times H\times W(C=2)$ Wave function ($u$)
NS (FNO [4]) Vorticity ($w$), Spatiotemporal Coordinates $H\times W\times T\times 4(T=33)$ Vorticity ($w\in[0,T)$)
NS (PDEBench [10]) Velocity ($v_{x},v_{y}$), pressure ($p$) $T\times C\times H\times W(T=15,C=3)$ Velocity ($v_{x},v_{y}$), pressure ($p$) at $T+1$
RD (PDEBench [10]) Activator ($u$), inhibitor ($v$) $T\times C\times H\times W(T=10,C=2)$ Activator ($u$), inhibitor ($v$) at $T+1$
@@ B.1 Distributions of Unlabeled PDE Data
In Table 2, for the purpose of OOD testing, we summarize the distribution of our unlabeled PDE data during pretraining, fine-tuning, and inference with in-context examples. Ranges of these physical parameters are inspired by [11]. During pretraining, we consider a wide distribution of unlabeled PDE data. When training (fine-tuning) our model, we consider in-distribution unlabeled PDE data. Finally, we test our similarity-based method that learns in-context examples in OOD settings.333 For Navier Stokes from PDEBench, we use the original data. We could not generate our own pretraining/finetuning data with different ranges of physics parameters, due to a possible mismatch of the provided configuration files and the version of Phiflow used in PDEBench (see GitHub issue at https://github.com/pdebench/PDEBench/issues/36). For Helmholtz OOD, we choose a narrow range of coefficients ([15, 20]) mainly because its solution is very sensitive to the wavenumber, and FNO’s performance significantly drops when we move to more extreme OOD settings.
Table 2: Ranges of physical parameters (integers) for unsupervised pretraining, training (fine-tuning), and out-of-distribution (OOD) inference.
@@ B.3 Training Hyperparameters
64$\times$64 64$\times$64 128$\times$128$\times$10 64$\times$64$\times$33 512$\times$512$\times$15 180$\times$180$\times$16 180$\times$120$\times$16 96$\times$45
@@
Rollouts N/A N/A 91 N/A 1 1 1 N/A
Appendix C Examples of Simulation Costs
@@ Appendix D Model Architectures
In Figure 7, We show visualizations of architectures described in Sec. 3.1.4. Specifically, the FNO has 67.1M parameters, and the Video-MAE has 23.4M. During pretraining, FNO costs 20 GPU hours, and Video-MAE costs 18 GPU hours. During fine-tuning, FNO costs 4 GPU hours, and Video-MAE costs 6 GPU hours.
@@ Appendix E Comparison with Contrastive Learning
Appendix F More Comparisons with Vision Pretrained Models
Beyond Figure 3(e) for the transformer on time-dependent PDE (Navier-Stokes), we further verify that pretraining on unsupervised PDE data makes FNO outperform its off-the-shelf vision-pretrained checkpoints. Specifically, we first pretrain the 2D FNO model on ImageNet [87], then fine-tune on the downstream PDE simulation data. As shown in Figure 9, vision-pretrained FNO performs worse during downstream fine-tuning on time-independent PDEs (including both Poisson and Helmholtz equations), confirming that domain-specific pretraining, even on unsupervised PDE data, is more beneficial than conventional checkpoints pretrained on unrelated domains like computer vision.
Figure 9: Pretraining neural operators on unlabeled PDE data improves its performance and data efficiency on Poisson (left), Helmholtz (right). “random init.”: models are trained from scratch with random initialization. “vision pretrained”: fine-tuning from the checkpoint pretrained on computer vision dataset ImageNet [87].
Appendix G Joint Pretraining Further Improves Performance
We also study if joint pretraining on unlabeled multiple PDE data can bring extra benefits. We combine all 46,080 unlabeled Poisson samples and all 46,080 unlabeled Helmholtz samples (see Table 3). We choose this setting because recent works on SciML foundation models [47, 86] also use all samples from each PDE for pretraining. We do zero-paddings for mismatched channels. From Figure 10, we can see that joint pretraining can further improve the performance of fine-tuning on different PDEs.
Figure 10: Joint unsupervised pretraining on multiple PDEs (green solid curve) further improves the data efficiency of neural operators when fine-tuning on Poisson (left), Helmholtz (middle), Reaction-Diffusion (right). “random init.”: models are trained from scratch with random initialization. “unsupervised”: models are pretrained on a single unsupervised PDE data. “unsupervised joint”: models are pretrained on a joint of multiple unsupervised PDE datasets. “NS”: Navier Stokes. “RD”: Reaction-Diffusion.
Appendix H Fine-tuning on Unseen PDEs is Challenging
We also try to fine-tune neural operators on unseen PDEs (i.e. PDEs different from pretraining). Mismatched channels are padded with zeros. We find this will lead to worse performance compared with models pretrained on the same PDE, as shown in Figure 11.
Figure 11: Fine-tuning FNO (pretrained on Poisson) on unseen samples from Helmholtz.
Appendix I Consistently Improved Generalization
@@
Appendix J More Ablation Studies
J.1 Magnitude of Perturbations during Pretraining.
@@
J.2 Ablation of the Number of Pretraining Samples.
@@
Appendix K Visualization of MAE Pretraining
@@
Appendix L Details and Visualizations of Real-World Data
L.1 ECMWF Reanalysis v5 (ERA5)
We utilize data from 2006 to 2015, with the snapshots taken every 6 hours and a spatial resolution of $360\times 360$. The total number of snapshots is 14600. We apply the mean-standard deviation normalization to the data and downsample the snapshots to a spatial resolution of $180\times 180$. We split 75% of the data for pretrain and 25% for finetune. For each split, we further separate 80% of the data for training, 10% for validation, and 10% for testing.
L.2 ScalarFlow
The original spatial resolution is $1062\times 600$. We crop the snapshots to $900\times 600$ to remove the padding and background. We remove the first 15 timeframes for each simulation to avoid the initial transient phase at the beginning of the smoke flow generation. We then downscale the snapshots to $180\times 120$ and apply mean-standard deviation normalization to the data. With a total of 70200 snapshots, we split 80% of the data for pretrain and 20% for finetune. For each split, we further separate 80% of the data for training, 10% for validation, and 10% for testing.
L.3 Airfoil
We split 80% of the training data for pretrain and 20% for finetune. The original spatial resolution of each sample is $128\times 128$. We crop the snapshots to $96\times 45$ to remove the background. We then apply min-max normalization channel-wise to the sample.
We show visualizations of real-world scientific data in Figure 15.
Figure 15: We show snapshot examples from ERA5 temperature [78] (a, b) and ScalarFlow [81] (c, d) at different temporal steps; and also an example of Airfoil mask, velocities, and pressure [75] (e-j).
Appendix M Benefits of In-context Examples
@@
We try to understand the benefit of leveraging in-context examples by decomposing the relative MSE error into “scale” and “shape.” “Scale” is the slope of a linear regression between targets and predictions (closer to 1 the better), indicating the alignment of the range of model outputs with targets. “Shape” is the normalized relative MSE (i.e., model outputs or targets are normalized by their own largest magnitude before MSE), indicating the alignment of scale-invariant spatial/temporal structures. We show results in Figure 16. We find that the benefit of in-context examples lies in that the scale of the model’s output keeps being better calibrated (“scale” being closer to 1) when adding more demos.
In numerical simulations or predictions of PDEs, there are settings where the scale or magnitude of solutions is more important than the exact shape/pattern:
  1. 1.
    Heat Transfer: In large-scale systems, the focus might be on overall temperature and extreme values. For instance, in evaluating a cooling system, the key concern might be the peak temperature rather than the detailed temperature distribution.
  2. 2.
    Fluid Dynamics: For applications like aerodynamics, the overall drag or lift force on an object is often more critical than capturing every detail of the flow pattern, such as in airfoil design.
  3. 3.
    Environmental Modeling: The concentration of pollutants at specific locations or total pollutant transport is often more crucial than the exact distribution, such as in groundwater flow studies.
@@
Figure 16: Benefits of in-context examples. To analyze the benefit of in-context examples for complicated PDE systems, we decompose the relative MSE error into “Scale” and “Shape”. “Scale” indicates the alignment of the range of model outputs with targets (closer to 1 the better), via the slope of a linear regression. “Shape” indicates the alignment of scale-invariant spatial/temporal structures via normalized relative MSE (i.e. model outputs or targets are normalized by their own largest magnitude before MSE). We find that the benefit of in-context examples lies in that the scale of the model’s output keeps being calibrated (red line being closer to 1) when adding more demos.
Appendix N Visualizations with In-Context Examples
We show visualizations of our similarity-based mining of in-context examples in Figure 17. In this visualization, we find that the range of numerical solutions (e.g., values in colorbars) predicted with in-context examples becomes closer to the target. Meanwhile, based on this visualization, we conclude that OOD generalization is challenging for neural operators because of: 1) Significantly different patterns of solutions under different physical parameters; 2) Different value ranges of solutions under different physical parameters.
Figure 17: Visualizations of mining in-context examples for FNO in OOD testing. Ranges of solutions predicted with in-context examples (min/max of each snapshot, reflected in colorbars) become closer to the target.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes] .
  4. Justification: All claims stated in the abstract and introduction are evaluated in our experiments (Sec. 4).
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes] .
  9. Justification: We discussed our limitations in a separate Section 6.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A] .
  14. Justification: This paper does not include theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes] .
  19. Justification: We include instructions for reproducibility in the Appendix, including choices of unlabeled PDE data (Appendix A), data generation and training hyperparameters B, and model structures D. We also attached our code in supplement.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes] .
  24. Justification: We attached our code in the supplement.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes] .
  29. Justification: We include our experimental settings in Appendix B.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes] .
  34. Justification: All our operator pretraining experiments are reproduced for three times.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes] .
  39. Justification: We discussed the compute resources we used in Appendix B.3.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes] .
  44. Justification: Our work does not involve any potential harm, societal impact, or violation.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A] .
  49. Justification: This work is about foundational research on data augmentations. It does not lead to any societal impact because: 1) our motivations, goals, and results are too distant from social applications to make any social impact; 2) we mainly use public data and benchmarks; and 3) we use public model architectures and follow previous works which have open access.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A] .
  54. Justification: The paper poses no such risks because all datasets studied in our work are from public benchmarks, and we have clearly cited original papers.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes] .
  59. Justification: All datasets studied in our work are from public benchmarks, and we have clearly cited original papers.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A] .
  64. Justification: The paper does not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A] .
  69. Justification: The paper does not involve crowdsourcing or research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A] .
  74. Justification: The paper does not involve any potential risk regarding research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.